# Short geodesics and multiplicities of eigenvalues of hyperbolic surfaces

### v1

@@ 1. Introduction
$\mathcal{N}_{\epsilon}(X_{g}):=\{\gamma;\ \text{$\gamma\subset X_{g}$ is a simple closed geodesic of length $<2\epsilon$}\},$
@@
    This was due to Letrouit-Machado [LM24] for this case. We remark here that Jiang-Tidor-Yao-Zhang-Zhao [JTY+21] proved the same sublinear upper bound for the second largest eigenvalue of connected $g$-vertex graphs with bounded maximum degree.
@@
    The linear upper bound was first obtained by Besson [Bes80].
@@
A key novel ingredient in proving Theorem 1.1 is to find a linear subspace $S$ of the eigenspace $E(\lambda_{1}(X_{g}))$ with respect to $\lambda_{1}(X_{g})$ such that the dimension $\dim(S)$ of the subspace $S$ is uniformly comparable to $\dim(E(\lambda_{1}(X_{g})))=m_{1}(\lambda_{1}(X_{g}))$, and the $L^{2}$-norm of each eigenfunction in $S$ on $X_{g}$ is uniformly comparable to its $L^{2}$-norm on the $\epsilon$-thick part of $X_{g}$. And then restricted to the thick part of $X_{g}$, we follow a strategy similar to that in the recent works [JTY+21, LM24] to complete the proof of Theorem 1.1.
@@
Figure 2. A closed hyperbolic surface $\mathcal{Y}_{g}$ of genus $g$ separated by $g+1$ geodesics into two pieces, where $\gamma_{i}\in\mathcal{N}_{\epsilon}(\mathcal{Y}_{g})$ and $\epsilon=o(1)$ as $g\to\infty$. So $m(\lambda_{1}(\mathcal{Y}_{g}))=o(g)$.
We remark here that for a family of closed hyperbolic surfaces $\{X_{g}^{t}\}_{t\in(0,1]}$ of genus $g$ that has $(3g-3)$ pairwise disjoint simple closed geodesics simultaneously pinching to zero as $t\to 0$, it is known from e.g., [SWY80] or [HW25, Proof of Theorem 21] that $\lambda_{2g-3}(X_{g}^{t})\to 0$ uniformly as $t\to 0$. However, both Theorem 1.1 and Theorem 1.2 cannot distinguish the first $(2g-3)$ nonzero eigenvalues of this particular family $\{X_{g}^{t}\}_{t\in(0,1]}$. To our knowledge, it is not even known whether the multiplicity $m(\lambda_{1}(X_{g}^{t}))$ is sublinear in $g$ when $t$ is close to 00.
@@
The first upper bound on the multiplicity of Laplacian eigenvalues on closed Riemannian surfaces was established by Cheng [Che76], who showed that for any closed Riemannian surface of genus $g$, the multiplicity of the $i$-th eigenvalue $\lambda_{i}$ satisfies
@@
This was later improved by Besson [Bes80], who obtained the first linear upper bound:
@@
Many refinements of Besson’s bound have since been obtained. Nadirashvili [Nad88] improved the bound for $g\geq 2$ to $14g+2i-$ 1; Sévennec [Sév02] improved it further for $g\geq 2$, $i=1$ to $32g+$ 3; and Bourque–Petri [FBP23] improved the bound for $i=1$ to $12g-$ 1 for hyperbolic surfaces of sufficiently large genus.
@@
where chromatic number $\mathrm{chr}(M_{g})$ is the supremum of the natural numbers $n$ for which the complete graph on $n$ vertices can be embedded in $M_{g}$. It is known from [RY68] of Ringel-Youngs that
@@
demonstrating that the maximal multiplicity grows at least on the order of $\sqrt{g}$. However, counterexamples to the conjecture are found for $g=10$ and $g=17$ by Bourque-Gruda Mediavilla-Petri-Pineault in [BGMPP23].
Nevertheless, for closed hyperbolic surfaces of high genus, it is quite interesting to determine the right growth rate of the maximal multiplicity of the first nonzero eigenvalue of the Laplacian. Is it $\sqrt{g}$ in the above conjecture of Colin de Verdière, or $\frac{g}{\log\log g}$ as shown above, or something else? Several sublinear upper bounds on the multiplicity of the first nonzero eigenvalue of the Laplacian on hyperbolic surfaces have been established under certain assumptions [LM24, FBP23].
@@
[…]
@@ 2. Thick-thin decomposition
is always embedded in $X_{g}$. We call it the standard collar of $\gamma$. Moreover, $\operatorname{\mathcal{C}}(\gamma,w(\gamma))$ is isometric to $[-w(\gamma),w(\gamma)]\times\mathbb{R}/\mathbb{Z}$ endowed with the hyperbolic metric
@@
Lemma 2.1 (Collar Lemma, [Bus92, Theorem 4.1.6]).
@@
(3) $0<\epsilon<\operatorname{arcsinh}\frac{1}{\sinh 2}\approx 0.272$
@@
$\mathcal{N}_{\epsilon}(X_{g})=\{\gamma;\ \text{$\gamma\subset X_{g}$ is a simple closed geodesic of length $<2\epsilon$}\}.$
@@ 3. Mass distribution of eigenfunctions
$-\Delta=\frac{\partial^{2}}{\partial\rho^{2}}+\tanh\rho\frac{\partial}{\partial\rho}+\frac{1}{\ell(\gamma)^{2}\cosh^{2}\!\rho}\frac{\partial^{2}}{\partial t^{2}}.$
@@
Then $\sqrt{\cosh\rho}\cdot\alpha_{j}$ and $\sqrt{\cosh\rho}\cdot\beta_{j}$ satisfy the following differential equation (one may refer to e.g., [Gam02, Section 4], [Mon15, Section 3] and [WZ25, Section 5] for more details):
@@
By plugging $u_{1}=\varphi_{j}\textit{ or }\psi_{j}$ and $u_{2}=\cosh\rho$ into [WZ25, Lemma 5.2], then we have
@@
$\varphi_{j,\gamma}^{1}(x)=\begin{cases}\varphi_{j}(\rho)\cos(2\pi jt),&x=(\rho,t)\in\operatorname{\mathcal{C}}(\gamma,w(\gamma));\\ 0,&x\notin\operatorname{\mathcal{C}}(\gamma,w(\gamma)).\end{cases}$
@@
$\varphi_{j,\gamma}^{2}(x)=\begin{cases}\varphi_{j}(\rho)\sin(2\pi jt),&x=(\rho,t)\in\operatorname{\mathcal{C}}(\gamma,w(\gamma));\\ 0,&x\notin\operatorname{\mathcal{C}}(\gamma,w(\gamma)).\end{cases}$
@@
Recall that for $10<\epsilon<$ 1,
$\mathcal{N}_{\epsilon}(X_{g})=\{\gamma;\ \text{$\gamma\subset X_{g}$ is a simple closed geodesic of length $<2\epsilon$}\}.$
@@
is a fixed constant, and $N_{\epsilon}(X_{g})=\#\mathcal{N}_{\epsilon}(X_{g})$. By the general upper bound of the first nonzero eigenvalue (e.g., [Che75, Corollary 2.3] or [WZ25, Proposition 9.1]), one has that as $g\to\infty$,
@@ 4.1. Rescaling the metric
We first rescale the hyperbolic metric on $X_{g}$ by $\epsilon^{-2}$ to obtain a new Riemannian surface $X_{g,\epsilon}$. Then the 2-dimensional simply connected space form of curvature $-\epsilon^{2}$, denoted by $\mathbb{H}_{\epsilon}$, is the universal cover of $X_{g,\epsilon}$. Let $\operatorname{dist}_{\epsilon}(\cdot,\cdot)$ denote the distance function on $X_{g,\epsilon}$, and for any $x\in X_{g,\epsilon}$, let $B_{\epsilon}(x,r)$ denote the geodesic ball of radius $r>0$ centered at $x$.
@@
(15) $K_{t,\epsilon}(x,y)=\sum_{\gamma\in\Gamma}k_{t,\epsilon}(\gamma\tilde{x},\tilde{y}),$
where $\tilde{x}$ and $\tilde{y}$ are lifts of $x$ and $y$, respectively. Since $k_{t,\epsilon}(\tilde{x},\tilde{y})$ depends only on $\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y})$, we also write it as $k_{t,\epsilon}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y}))$. Next, we present a standard estimate for the heat kernel on $X_{g,\epsilon}$ which will be used in the subsequent analysis.
@@
$k_{t,\epsilon}(\tilde{x},\tilde{y})=k_{t,\epsilon}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y}))=\epsilon^{2}k_{\epsilon^{2}t,1}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y})),$
it follows from [Ber16, Lemma 3.13] that for all $t\geq 1$,
@@
(17) #$\#\{\gamma\in\Gamma|\ m\leq\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\gamma\tilde{x},\tilde{y})<m+1\}\leq\frac{\operatorname{Area}\big{(}B_{\mathbb{H}_{\epsilon}}(m+1+1/2)\big{)}}{\operatorname{Area}\big{(}B_{\mathbb{H}_{\epsilon}}(1/2)\big{)}}$
@@
(18) $r_{1}=c\log\log\Big(\frac{Kg}{N_{\epsilon}+1}\Big{missing}),\qquad r_{2}=c\log\Big(\frac{Kg}{N_{\epsilon}+1}\Big{missing}),$
@@ 4.2. Reduce the problem to the thick part
Following the ideas of [LM24, JTY+21], we will consider the multiplicity of the eigenvalue $e^{-r_{1}\lambda_{1,\epsilon}}$ in the spectrum of the operator $Pe^{-r_{1}\Delta_{\epsilon}}P$. First, we have
@@ 4.3.1. Estimate on $Y_{1}$
Since the operator norms of $P$, $e^{-r_{1}\Delta_{\epsilon}}$ and $e^{-(r_{1}-1)\Delta_{\epsilon}}$ are all no more than $1$, we have
@@
[…]
@@ 4.3.2. Estimate on $Y_{2}$
We omit the subscript $L^{2}(X_{g,\epsilon})$ in the following for simplicity. Following the approach in [LM24, Section 3 and Section 4] of Letrouit-Machado, we give an upper bound for
@@
These facts allow us to repeat the arguments in the proofs of [LM24, Lemma 3.1, 3.4, 3.8], taking $\rho=1$ and $b=-1$ in the initial assumptions of [LM24, Section 3], to derive the following three special cases of those lemmas.
Lemma 4.10 ([LM24], Lemma 3.1).
@@
Following [LM24], we denote by $\varphi_{x}\in L^{2}(X_{g,\epsilon})$ a function that attains the following supremum
@@
[…]
@@
[…]
@@
$C_{5}\exp(r_{1}-\frac{C^{2}r_{1}}{64})\leq C_{5}\exp\big(-hr_{1}-3r_{1}\big{missing})\leq\frac{1}{4}C_{4}\exp(-hr_{1})e^{-2r_{1}\lambda_{1,\epsilon}}.$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
(36) $X^{\mathrm{thin}}_{g}$ $:=\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma)-\frac{1}{\delta}),$
$X^{\mathrm{thick}}_{g}$ $:=\overline{X\setminus X^{\mathrm{thin}}_{g}}=\mathcal{A}_{1}\sqcup\mathcal{A}_{2}\sqcup\cdots\sqcup\mathcal{A}_{I_{\epsilon}},\quad I_{\epsilon}\geq 1.$
@@
$\operatorname{inj}(x)>\epsilon\text{\quad and\quad}B(x,\epsilon)\subset\mathcal{A}_{k}\cup\mathcal{S}_{k}.$
@@
Otherwise, $\ell(\alpha)\geq 2$, while by (5), $\ell(\alpha_{pq})<2$, which leads to a contradiction since we assume that $\alpha$ is shortest. Thus, $d\leq 1/\delta+1$, and by (2) we have
@@
Suppose $p,q\in\mathcal{A}_{k}$, and let $\alpha$ be a shortest geodesic segment joining $p$ and $q$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$. Then
|$\absolutevalue{f(p)-f(q)}^{2}\leq c(\epsilon)\cdot\lambda\cdot\ell(\alpha)$
@@
The proof is essentially the same as the proof of [WX22, Proposition 12]. Let $\alpha:[0,\ell]\to X_{g}$ be parameterized by the arc-length parameter such that $\alpha(0)=p$ and $\alpha(\ell)=q$, where $\ell=\ell(\alpha)$. For any $x\in\alpha$, by Lemma 5.1 we have $\operatorname{inj}(x)>\epsilon$. Thus, by standard Sobolev embedding theorem [Tay23],
$\norm{\nabla f}_{L^{\infty}(B(x,\epsilon/2))}\leq\tilde{c}(\epsilon)\sum_{j=0}^{N}\norm{\Delta^{j}(\mathop{}\!\mathrm{d}f)}_{L^{2}(B(x,\epsilon))}\leq\frac{4}{3}\tilde{c}(\epsilon)\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(B(x,\epsilon))},$

### v2

@@ 1. Introduction
[…]
@@
    This was due to Letrouit–Machado [LM2024] for this case. We remark here that Jiang–Tidor–Yao–Zhang–Zhao [JTYZZ2021] proved the same sublinear upper bound for the second largest eigenvalue of connected $g$-vertex graphs with bounded maximum degree.
@@
    The linear upper bound was first obtained by Besson [Besson1980].
@@
A key novel ingredient in proving Theorem 1.1 is to find a linear subspace $S$ of the eigenspace $E(\lambda_{1}(X_{g}))$ with respect to $\lambda_{1}(X_{g})$ such that the dimension $\dim(S)$ of the subspace $S$ is uniformly comparable to $\dim(E(\lambda_{1}(X_{g})))=m_{1}(\lambda_{1}(X_{g}))$, and the $L^{2}$-norm of each eigenfunction in $S$ on $X_{g}$ is uniformly comparable to its $L^{2}$-norm on the $\epsilon$-thick part of $X_{g}$. And then restricted to the thick part of $X_{g}$, we follow a strategy similar to that in the recent works [JTYZZ2021, LM2024] to complete the proof of Theorem 1.1.
@@
Figure 2. A closed hyperbolic surface $\mathcal{Y}_{g}$ of genus $g$ separated by $g+1$ geodesics into two pieces, where $\gamma_{i}\in\mathcal{N}_{\epsilon}(\mathcal{Y}_{g})$ and for each $i$, $\ell(\gamma_{i})\to 0$ as $g\to\infty$. So $m(\lambda_{1}(\mathcal{Y}_{g}))=o(g)$.
We remark here that for a family of closed hyperbolic surfaces $\{X_{g}^{t}\}_{t\in(0,1]}$ of genus $g$ that has (3g3)(3g-3) pairwise disjoint simple closed geodesics simultaneously pinching to zero as $t\to 0$, it is known from e.g., [SWY80] or [HW25, Proof of Theorem 21] that $\lambda_{2g-3}(X_{g}^{t})\to 0$ uniformly as $t\to 0$. However, both Theorem 1.1 and Theorem 1.2 cannot distinguish the first (2g3)(2g-3) nonzero eigenvalues of this particular family $\{X_{g}^{t}\}_{t\in(0,1]}$. To our knowledge, it is not even known whether the multiplicity $m(\lambda_{1}(X_{g}^{t}))$ is sublinear in $g$ when $t$ is close to 0.
@@
The first upper bound on the multiplicity of Laplacian eigenvalues $\{\lambda_{i}\}$ on closed Riemannian surfaces was established by Cheng [Cheng1976], who showed that for any closed Riemannian surface of genus $g$, the multiplicity of the $i$-th eigenvalue $\lambda_{i}$ satisfies
@@
This was later improved by Besson [Besson1980], establishing the first linear upper bound:
@@
Many refinements of Besson’s bound have since been obtained. Nadirashvili [Nadirashvili1988] improved the bound for $g\geq 2$ to 4g+2i14g+2i-1; Sévennec [Sevennec2002] improved it further for $g\geq 2$, $i=1$ to $2g+3$; and Bourque–Petri [BP2023] improved the bound for $i=1$ to 2g12g-1 for hyperbolic surfaces of sufficiently large genus.
Several sublinear upper bounds on the multiplicity of the first nonzero Laplacian eigenvalue on hyperbolic surfaces have been established under certain assumptions. As introduced above, if the surface has injectivity radius $\geq\epsilon>0$, Letrouit–Machado [LM2024] proved that
$m(\lambda_{1})\ll_{\epsilon}\frac{g}{\log\log g}.$
Gilmore–Le Masson–Sahlsten–Thomas [GLeMST21] proved that if the surface has injectivity radius $\geq\frac{1}{(\log g)^{\alpha}}$, and at each point there is at most one geodesic loop of length less than $c\log g$ based there, then
$m(\lambda)$ $\ll_{c,\lambda}\frac{g}{(\log g)^{1-\alpha}}\text{\quad for\ }\lambda\in\Big(\frac{1}{4},+\infty\Big),$
$\text{and\quad}m(\lambda)$ $\ll_{c}g^{1-2c\sqrt{\epsilon}}(\log g)^{2\alpha}\text{\quad for\ }\lambda\in\Big(0,\frac{1}{4}-\epsilon\Big).$
Monk [Monk22] proved that if the surface has injectivity radius $\geq g^{-\frac{1}{24}}(\log g)^{\frac{9}{16}}$ and the area of the points where the injectivity radius $\frac{1}{6}\log g$ is $O(g^{\frac{2}{3}})$, then
$m(\lambda)\ll\sqrt{1+\lambda}\frac{g}{\sqrt{\log g}}.$
We emphasize here that our upper bound in Theorem 1.1 and Theorem 1.2 do not require any lower bound of the injectivity radius, instead we need the upper bounds of the number of short closed geodesics or components of the thick part of the surface. Bourque–Petri [BP2023] proved that, for every $p_{1},p_{2}\in(j_{0},\pi]$, where $j_{0}$ is the first positive zero of the Bessel function $J_{0}$,
$m(\lambda_{1})\ll_{p_{1},p_{2}}\frac{g}{(\log g)^{3}},\text{\quad when\ $\lambda_{1}\in\bigg[\frac{1}{4}+\Big(\frac{p_{1}}{\log g}\Big)^{2},\frac{1}{4}+\Big(\frac{p_{2}}{\log g}\Big)^{2}\bigg]$}.$
@@
where the chromatic number $\mathrm{chr}(M_{g})$ is the supremum of the natural numbers $n$ for which the complete graph on $n$ vertices can be embedded in $M_{g}$. It is known from [RY68] of Ringel–Youngs that
@@
demonstrating that the maximal multiplicity grows at least on the order of $\sqrt{g}$. However, counterexamples to the conjecture are found for $g=10$ and $g=17$ by Bourque–Gruda-Mediavilla–Petri–Pineault in [BGMPP23].
Nevertheless, for closed hyperbolic surfaces of high genus, it is quite interesting to determine the right growth rate of the maximal multiplicity of the first nonzero eigenvalue of the Laplacian. Is it $\sqrt{g}$ in the above conjecture of Colin de Verdière, or $\frac{g}{\log\log g}$ as shown above, or something else?
@@
For any two positive functions $f(g)$ and $h(g)$, we say $f=o(h)$ if $\lim\limits_{g\to\infty}\frac{f(g)}{h(g)}=0$; we say $f\ll h$ if there exists a uniform constant $C>0$ such that $f\leq Ch$. We use $\lfloor\cdot\rfloor$ to denote the greatest integer function.
@@ 2. Thick-thin decomposition
is always embedded in $X_{g}$. We call it the standard collar of $\gamma$. Moreover, $\operatorname{\mathcal{C}}(\gamma,w(\gamma))$ is isometric to $w(\gamma),w(\gamma)]\times\mathbb{R}/\mathbb{Z}$ endowed with the hyperbolic metric
@@
Lemma 2.1 (Collar Lemma, [Buser1992, Theorem 4.1.6]).
@@
(3) $0<\epsilon<\operatorname{arcsinh}\frac{1}{\sinh 2}\approx 0.272.$
@@
[…]
@@ 3. Mass distribution of eigenfunctions
$\Delta=\frac{\partial^{2}}{\partial\rho^{2}}+\tanh\rho\frac{\partial}{\partial\rho}+\frac{1}{\ell(\gamma)^{2}\cosh^{2}\!\rho}\frac{\partial^{2}}{\partial t^{2}}.$
@@
Then $\sqrt{\cosh\rho}\cdot\alpha_{j}$ and $\sqrt{\cosh\rho}\cdot\beta_{j}$ satisfy the following differential equation (one may refer to e.g., [Gamburd2002, Section 4], [Mondal2015, Section 3] and [WZ2025, Section 5] for more details):
@@
By plugging $u_{1}=\varphi_{j}\textit{ or }\psi_{j}$ and $u_{2}=\cosh\rho$ into [WZ2025, Lemma 5.2], then we have
@@
[…]
@@
[…]
@@
Recall that for $0<\epsilon<1$,
[…]
@@
is a fixed constant, and $N_{\epsilon}(X_{g})=\#\mathcal{N}_{\epsilon}(X_{g})$. By the general upper bound of the first nonzero eigenvalue (e.g., [Cheng1975, Corollary 2.3] or [WZ2025, Proposition 9.1]), one has that as $g\to\infty$,
@@ 4.1. Rescaling the metric
We first rescale the hyperbolic metric on $X_{g}$ by $\epsilon^{-2}$ to obtain a new Riemannian surface $X_{g,\epsilon}$. Then the 2-dimensional simply connected space form of curvature $\epsilon^{2}$, denoted by $\mathbb{H}_{\epsilon}$, is the universal cover of $X_{g,\epsilon}$. Let $\operatorname{dist}_{\epsilon}(\cdot,\cdot)$ denote the distance function on $X_{g,\epsilon}$, and for any $x\in X_{g,\epsilon}$, let $B_{\epsilon}(x,r)$ denote the geodesic ball of radius $r>0$ centered at $x$.
@@
[…]
@@
$),k_{t,\epsilon}(\tilde{x},\tilde{y})=k_{t,\epsilon}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y}))=\epsilon^{2}k_{\epsilon^{2}t,1}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y}$)),
it follows from [Bergeron2016, Lemma 3.13] that for all $t\geq 1$,
@@
(17) $\#\{\gamma\in\Gamma|\ m\leq\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\gamma\tilde{x},\tilde{y})<m+1\}\leq\frac{\operatorname{Area}\big(B_{\mathbb{H}_{\epsilon}}(m+1+1/2)\big)}{\operatorname{Area}\big(B_{\mathbb{H}_{\epsilon}}(1/2)\big)}$
@@
[…]
@@ 4.2. Reduce the problem to the thick part
Following the ideas of [LM2024, JTYZZ2021], we will consider the multiplicity of the eigenvalue $e^{-r_{1}\lambda_{1,\epsilon}}$ in the spectrum of the operator $Pe^{-r_{1}\Delta_{\epsilon}}P$. First, we have
@@ 4.3.1. Estimate on $Y_{1}$
Since the operator norms of $P$, $e^{-r_{1}\Delta_{\epsilon}}$ and $e^{-(r_{1}-1)\Delta_{\epsilon}}$ are all no more than 11, we have
@@
$\leq{}$ $\|K_{1,\epsilon}(x,\cdot)\|_{L^{2}(X_{g,\epsilon})}+\|\psi_{k_{x}}\|_{L^{\infty}(X_{g,\epsilon})}\cdot\|e^{-\Delta_{\epsilon}}\psi_{k_{x}}\|_{L^{2}(X_{g,\epsilon})}$
$\leq{}$ $\|K_{1,\epsilon}(x,\cdot)\|_{L^{\infty}(X_{g,\epsilon})}^{\frac{1}{2}}\cdot\|K_{1,\epsilon}(x,\cdot)\|_{L^{1}(X_{g,\epsilon})}^{\frac{1}{2}}+\frac{2}{\sqrt{\pi}}$
$\leq{}$ $\sqrt{C}_{1}+\frac{2}{\sqrt{\pi}}.$
@@ 4.3.2. Estimate on $Y_{2}$
We omit the subscript $L^{2}(X_{g,\epsilon})$ in the following for simplicity. Following the approach in [LM2024, Section 3–4] of Letrouit–Machado, we give an upper bound for
@@
These facts allow us to repeat the arguments in the proofs of [LM2024, Lemma 3.1, 3.4, 3.8], taking $\rho=1$ and b=1b=-1 in the initial assumptions of [LM2024, Section 3], to derive the following three special cases of those lemmas.
Lemma 4.10 ([LM2024], Lemma 3.1).
@@
Following [LM2024], we denote by $\varphi_{x}\in L^{2}(X_{g,\epsilon})$ a function that attains the following supremum
@@
$\ll{}$ $\int_{Y_{2}}\norm{(P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P)^{n}\varphi_{x}}^{2}+\exp(-\frac{C^{\prime 2}r_{2}}{16})\mathrm{d}A_{x}$
$\ll{}$ $\left(\int_{Y_{2}\setminus Y_{3}}+\int_{Y_{3}}\right)\norm{(P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P)^{n}\varphi_{x}}^{2}\mathrm{d}A_{x}+\epsilon^{-2}g\exp(-\frac{C^{\prime 2}r_{2}}{16})$
$\ll{}$ $\int_{Y_{2}\setminus Y_{3}}\|(P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P)^{n}\varphi_{x}\|^{2}\mathrm{d}A_{x}+\exp(3C^{\prime}r_{2})+\epsilon^{-2}g\exp(-\frac{C^{\prime 2}r_{2}}{16}),$
@@
$\leq{}$ $\big(1-C_{4}\exp(-hr_{1})\big)\cdot\big(e^{-2r_{1}\lambda_{1,\epsilon}}+C_{3}\exp(-C^{\prime}r_{2})\big)$
$\leq{}$ $\big(1-\frac{1}{2}C_{4}\exp(-hr_{1})\big)\cdot e^{-2r_{1}\lambda_{1,\epsilon}}.$
@@
$C_{5}\exp(r_{1}-\frac{C^{2}r_{1}}{64})\leq C_{5}\exp\big(-hr_{1}-3r_{1}\big.)\leq\frac{1}{4}C_{4}\exp(-hr_{1})e^{-2r_{1}\lambda_{1,\epsilon}}.$
@@
$={}$ $\int_{Y_{2}\setminus Y_{3}}\|P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P\varphi_{x}\|^{2{n}}\mathrm{d}A_{x}$
$\ll{}$ $\epsilon^{-2}g\cdot\Big(1-\frac{1}{4}C_{4}\exp(-hr_{1})\Big)^{n}e^{-2r_{1}n\lambda_{1,\epsilon}}$
$\leq{}$ $\epsilon^{-2}g\cdot\exp(-\frac{1}{4}C_{4}\exp(-hr_{1})n)e^{-2r_{1}n\lambda_{1,\epsilon}},$
@@
$={}$ $\epsilon^{-2}(N_{\epsilon}+1)\Big(\frac{Kg}{N_{\epsilon}+1}\Big)^{(4+3C^{\prime})c}\ll\epsilon^{-2}(g(N_{\epsilon}+1))^{\frac{1}{2}}.$
@@
$={}$ $\epsilon^{-2}g\Big(\frac{Kg}{N_{\epsilon}+1}\Big)^{\big(4-\frac{C^{\prime 2}}{16}\big)c}\leq\epsilon^{-2}g\Big(\frac{Kg}{N_{\epsilon}+1}\Big)^{4c-1}\ll\epsilon^{-2}(g(N_{\epsilon}+1))^{\frac{1}{2}}.$
@@
$\leq{}$ $\epsilon^{-2}g\cdot\exp(2r_{1}-\frac{1}{4}C_{4}\exp(-\frac{1}{4c}r_{1})n)$
$\ll{}$ $\epsilon^{-2}g\cdot\exp(-\frac{1}{5}C_{4}\Big(\log\frac{Kg}{N_{\epsilon}+1}\Big)^{\frac{2}{3}})$
$\ll{}$ $\epsilon^{-2}\frac{g}{\log\frac{Kg}{N_{\epsilon}+1}},$
@@
(36) $X^{\mathrm{thin}}_{g}$ $=\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma)-\frac{1}{\delta}),$
$X^{\mathrm{thick}}_{g}$ $=\overline{X\setminus X^{\mathrm{thin}}_{g}}=\mathcal{A}_{1}\sqcup\mathcal{A}_{2}\sqcup\cdots\sqcup\mathcal{A}_{I_{\epsilon}},\quad I_{\epsilon}\geq 1.$
@@
$\operatorname{inj}(x)>\epsilon\text{\qquad and\qquad}B(x,\epsilon)\subset\mathcal{A}_{k}\cup\mathcal{S}_{k}.$
@@
Otherwise, $\ell(\alpha)\geq 2$, while by (5), $\ell(\alpha_{pq})<2$, which leads to a contradiction since we assume that $\alpha$ is the shortest. Thus, $d\leq 1/\delta+1$, and by (2) we have
@@
Suppose $p\in\mathcal{A}_{k}$, $q\in\mathcal{A}_{k}\cup\mathcal{S}_{k}$ and let $\alpha$ be a shortest geodesic segment joining $p$ and $q$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$. Then
[…]
@@
The proof is essentially the same as the proof of [WX2022b, Proposition 12]. Let $\alpha:[0,\ell]\to X_{g}$, $\ell=\ell(\alpha)$ be the arc-length parameterization such that $\alpha(0)=p$ and $\alpha(\ell)=q$. For any $x\in\alpha$, by an argument similar to Lemma 5.1 we may assume $\operatorname{inj}(x)>\epsilon/2$. Thus, by standard Sobolev embedding theorem [Taylor2023],
$\norm{\nabla f}_{L^{\infty}(B(x,\epsilon/4))}\leq\tilde{c}(\epsilon)\sum_{j=0}^{N}\norm{\Delta^{j}(\mathop{}\!\mathrm{d}f)}_{L^{2}(B(x,\epsilon/2))}\leq\frac{4}{3}\tilde{c}(\epsilon)\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(B(x,\epsilon/2))},$